OpenStudy (anonymous):

What is the value of x? A. 5 B. 7 C. 10 D. 11

OpenStudy (anonymous):

x/12=55/60

OpenStudy (anonymous):

@tanya123 @timo86m @TSwizzle

geerky42 (geerky42):

I would start by multiply both sides of this equation by 12.

undeadknight26 (undeadknight26):

I would do this: 55/60 = .91 .91 x 12 = 11.

geerky42 (geerky42):

basically the same thing

OpenStudy (anonymous):

so... what is it

OpenStudy (anonymous):

i think you cross multiply

OpenStudy (anonymous):

whats te answer

geerky42 (geerky42):

can't you do what I and undedknight did? Just multiply both side by 12 to have x isolated. then simplify the right side. then there's your answer.

geerky42 (geerky42):

\[\dfrac{x}{12} = \dfrac{55}{60} \\ 12 \times\dfrac{x}{12} = \dfrac{55}{60} \times 12 \\ x = \dfrac{55}{60} \times 12\]

geerky42 (geerky42):

use calculator.

OpenStudy (anonymous):

@Jadeishere

OpenStudy (anonymous):

@Tazmaniadevil @Warriorz13

geerky42 (geerky42):

We are here to help you with math problems, not give you answer.

OpenStudy (anonymous):

@whpalmer4 @WinryRockbell @wagnermar @Waraiders00 @e.mccormick

OpenStudy (anonymous):

well I wisk you werent here

geerky42 (geerky42):

Do you have calculator you can use...?

OpenStudy (warriorz13):

I think the answer is 11

OpenStudy (warriorz13):

And you can just use google calculator =)

geerky42 (geerky42):

well, use calculator and finish this problem, then you don't have to deal with me and this problem.

OpenStudy (anonymous):

it is thanks f uck you all god dammit

OpenStudy (warriorz13):

>:( LEAVE @pinksparkles ALONE!!!

geerky42 (geerky42):

good luck with tests where there will be no access to internet...

OpenStudy (jadeishere):

....

OpenStudy (warriorz13):

Your right you are geerky you freaking jerk

OpenStudy (jadeishere):

x = 55/60 * 12 so.. 0.91 * 12 is x.. Not sure if you need that Or is it a ratio? Just curious... do I need to be here?

OpenStudy (anonymous):

no i got it thanks

OpenStudy (jadeishere):

Kay.. Sorry. But was it a ratio? That would change the way the answer came out..

OpenStudy (anonymous):

idk i failed the test

OpenStudy (anonymous):

You have completed the Lesson Checkpoint. Review the correct answer to each problem. If you did not master the Checkpoint, refer to the Extra Practice section of the Student Guide. 100% Assessment Summary Score: 5/5 points Status: Mastered View all scores (2) Retest Information You may retake this assessment only one more time. If you retake this assessment, only your new score — and not your previous scores — will count. Also note that the retest may have different questions from the original assessment. ReviewRetake